OpenStudy (blackops2luvr):

What is the length, in units, of the hypotenuse of a right triangle if each of the two legs is 2 units? 4 units 2 units 6sqrt units 8sqrt units

OpenStudy (blackops2luvr):

@iGreen

jimthompson5910 (jim_thompson5910):



OpenStudy (blackops2luvr):

4

jimthompson5910 (jim_thompson5910):

Use the pythagorean theorem to solve for c a^2 + b^2 = c^2 2^2 + 2^2 = c^2 ... ... ... c = ??